OpenStudy (anonymous):

ax+b < 0 is strict inequality ax^2 +bx + c > 0 is slack inequality, what does this mean?

OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 14 x + 3y = 9

OpenStudy (anonymous):

I would have said both were strict. (You can introduce a slack variable to an inequality for some purposes).

OpenStudy (anonymous):

@Zahabaha x-2y=14 x+3y=9 taking the first equation x-2y=14 x=14+2y putting this value in the second equation x+3y=9 (14+2y)+3y=9 14+5y=9 5y=9-14 5y=-5 dividing both sides by 5, we get y=-1 putting the value of y in x x=14+2y x=14-2 x=12